4x + 3y = –10
9x + 2y = 6

Answer by Boreal(15235)

4x + 3y = –10
9x + 2y = 6
------------------
8x+6y=-20, multiplying all by 2
-27x-6y=-18, multiplying all by -3
add them, and the y disappear
-19x=-38
x=2
go to the first
4(2)+3y=-10
8+3y=-10
3y=-18
y=-6
(2,-6)ANSWER
check in second equation
9(2)+2(-6)=6, and that checks